Find the present value of the given annuity due.

$800 paid at the beginning of each six-month period for six years at the rate 7% compounded semiannually.
Answer #1

Answer-

Amount = PMT = $ 800
Number of periods = N = 6 X 2 = 12 [ Since it is paid every six months for six years]
Interest rate = I/Y = 7 % compounded semiannually = 7 % / 2 = 3.5 %
Setting the financial calculator on beginning mode
PMT = - $ 800
N = 12
I/Y = 3.5
FV = 0
PV = $ 8001.24

Therefore the Present value of annuity due = $ 8001.24